The heat of fusion per gram of p-dibromobenzene, C4H4Br2, is 86.10 joules, the melting point is 86°c. Calculate the ideal solubility at 25°c
1
Expert's answer
2020-02-25T04:51:14-0500

Ideal solubility we can defined as the solubility of a given ideal solute in a given ideal solvent.

Here ,it is given that heat of fusion per gram for p- dibromobenzene, C6H4Br2 is 86.10 Joules

Tm= melting point of p-dibromobenzene is 86 0C

It is given the condition of ideal solubility as 25 0C

Using,Yalkowsky approximation for 250C

Log (s) = -0.01 ( Tm-25)

where s represents the ideal solubility

Log (s) = -0.01 (86-25)

Log(s) = -0.01"\times61"

Log(s)= -0.61

s= antilog (-0.61)

s= 2.454"\times" 10-1

This is the ideal solubility